If x > 0, show that there exists n ∈ N such that 1/n < x.


1
Expert's answer
2021-01-19T03:42:04-0500

Here we have "(x\in\R)>0" . Which implies "\frac{1}{x}>0" .


Now, using Archimedean property, we have "n\in \N" such that,



"0<\frac{1}{x}<n\\~\\
\Rightarrow n>\frac{1}{x}\\~\\
\Rightarrow x>\frac{1}{n}"

Hence, "\exists n\in \N" such that, "x>\frac{1}{n}"